OpenStudy (anonymous):

Solve by the substitution method. x+3y=12 8x+2y=-36

OpenStudy (haseeb96):

x+3y=12 ...........eq#1 8x+2y=-36..........eq#2 take eq#1, and separate x x=12-3y ........eq#3 put eq#3 into EQ#2 8(12-3y)+2y=-36 96-24y+2y=-36 96+36=22y y=132\22 y=6 now put the value of y into EQ # 3 x=12-3(6) x=12-18 x=-6 so the solution set is { (-6 , 6 ) } this is the correct answer